OpenStudy (anonymous):

Jade says she knows that 1/4 divided by 4= 1/16 because 1/16 *4 = 1/4. Is she correct? Why or why not? Medal to the best helper

OpenStudy (texaschic101):

1/4 divided by 4 = 1/4 * 1/4 = 1/16 1/16 * 4 = 4/16 which reduces to 1/4 correct. I don't really know how to answer why....it just does...lol

OpenStudy (texaschic101):

lol....I tried :(

OpenStudy (texaschic101):

ha...I found something. You can check your division with your multiplication. read here.. http://go.hrw.com/resources/go_sc/hst/HSTMW061.PDF thats how you know its correct